OpenStudy (anonymous):

can the set of lengths be the side lengths of a right triangle? 7 cm, 24 cm, 25cm.

OpenStudy (shamil98):

Test it out. a^2 + b^2 = c^2 7^2 + 24^2 = 25^2 if it is true, then it can be, if not it is false.